Ray and Amir each bought a used car for the same price. Each needed repairs. The repairs on Ray's car cost an additional $800 and on Amir's, $200. Amir noted, "Ray, my total cost was really only 5/6 of your total cost." How much did each pay for his used car?
Answer by mananth(16949)

let the price of their cars be $X
Ray cost = x+800
Amir's cost = x+200
x+200= 5/6 ( x+800)
6(x+200)= 5(x+800)
6x+1200 = 5x+4000
x= 4000-1200
x= $2800 price of car